OpenStudy (anonymous):

passing through (2, square root 30 / 3), distance between directrices 24 square root 7 / 7. find the equation of ellipse, center at origin.

OpenStudy (anonymous):

have you tried it?

OpenStudy (anonymous):

yes but i got stuck up when it comes to distance between directrices..

OpenStudy (anonymous):



OpenStudy (anonymous):

can you identify the directrices?

OpenStudy (anonymous):

sorry but no idea about directrices.. no, i mean its a/e and e is.. a/c. am i right?

OpenStudy (anonymous):

e is eccentricity.and yes directrices are at x=a/e and x=-a/e. so what's the distance between them?

OpenStudy (anonymous):

12 square root 7 / 7?

OpenStudy (anonymous):

no, i mean dist. btween them is 2a/e .

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

i dont understand.. how to get e.. no idea about getting a. im sorry

OpenStudy (anonymous):

i'll tell ,first you tell you understand how it is 2a/e?

OpenStudy (anonymous):

yes..its the distance between directrices

OpenStudy (anonymous):

\[2\frac{a}{e}=24 \frac{\sqrt{7}}{7}\] \[\frac{a}{e}=12 \frac{\sqrt{7}}{7}\] \[e= \frac{{7a}}{12\sqrt{7}}\]

OpenStudy (anonymous):

yes im following..

OpenStudy (anonymous):

also it passes through 2,root(30)/3 so,\[\frac{2^{2}}{a^{2}}+\frac{(\sqrt{30}/3)^{2}}{b^{2}}=1\]

OpenStudy (anonymous):

clear till now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

let, A=1/a^2 and B= 1/b^2 so our eq. becomes \[4A+\frac{30B}{9}=1\] clear?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

also \[e^2 = \frac{49a^2}{144*7}\] \[e^2 = \frac{7a^2}{144}\] \[e^2 = \frac{7}{144A}\] clear?

OpenStudy (anonymous):

yes yes

OpenStudy (anonymous):

Now... we have a relation ,you must remember \[e^2=1-\frac{b^2}{a^2}\] so \[e^2=1-\frac{A}{B}\] equating values of e^2. \[\frac{7}{144A}=1-\frac{A}{B}\] \[\frac{7}{144A}=\frac{B-A}{B}\] \[{7B}=144A(B-A)\] also we have a relation \[4A+\frac{30B}{9}=1\] solving these 2 relation you'll get A and B. clear till now??

OpenStudy (anonymous):

ill just equate the two relations?

OpenStudy (anonymous):

the first relation is the 4A + 30B/9 = 1 and?

OpenStudy (anonymous):

that 7B=144A(B-A)

OpenStudy (anonymous):

use substitution method.

OpenStudy (anonymous):

thank you so much ;))

OpenStudy (anonymous):

lifesaver

OpenStudy (anonymous):

and wait.. there were two equations in this question.. it was stated in the book

OpenStudy (anonymous):

i mean two answers

OpenStudy (anonymous):

yes ther are 2 answers. tell what values of A and B are you getting

OpenStudy (anonymous):

you'll get A=7/88,B=9/44 A=1/12,B=1/5

OpenStudy (anonymous):

yes right!!

OpenStudy (anonymous):

so the eq. of ellipse is: \[\frac{x^{2}}{a^{2}}+\frac{y^{2}}{b^{2}}=1\] \[Ax^2+By^2=1\]

OpenStudy (anonymous):

THANK YOU SO MUCH!! :)